An investment of $10,000 earned a net income of $345 in 1 year. Part of the investment was in bonds and earned income at a rate of 8%. The rest of the investment was in stocks and lost money at the rate of 5%. How much was invested in bonds, and how much was invested in stocks?
Answer by checkley77(12844)

.08x-.05(10,000-x)=345
.08x-500+.05x=345
.13x=345+500
.13x=845
x=845/.13
x=6,500 amount invested @ 8%.
10,000-6,500=3,500 amount invested @ 5%.
Proof:
.08*6,500-.05*3,500=345
520-175=345
345=345